SO, I am having trouble doing my math homework and cannot figure out the answer to this question. Here it is "7/11% of a quantity is equal to what fraction of the quantity" Please explain this so I could learn how to do this properly. Thanks.
Answer by Alan3354(69443)

"7/11% of a quantity is equal to what fraction of the quantity"
--------------
Per cent means per 100
1% = 1/100, etc.
----
7/11% = (7/11)/100
= 7/1100